# Exploring the effect of different cosmologies on the Epoch of Reionization 21-cm signal with Polar

### v1

@@ Abstract
A detection of the 21-cm signal power spectrum from the Epoch of Reionization is imminent, thanks to consistent advancements from telescopes such as LOFAR, MWA, and HERA, along with the development of SKA. In light of this progress, it is crucial to expand the parameter space of simulations used to infer astrophysical properties from this signal. In this work, we explore the role of cosmological parameters such as the Hubble constant $H_{0}$ and the matter clustering amplitude $\sigma_{8}$, whose values as provided by measurements at different redshifts are in tension. We run N-body simulations using Gadget-4, and post-process them with the reionization simulation code Polar, that uses L-Galaxies to include galaxy formation and evolution properties and Grizzly to execute 1-D radiative transfer of ionizing photons in the intergalactic medium (IGM). We compare our results with the latest JWST observations and explore which astrophysical properties for different cosmologies are necessary to match the observed UV luminosity functions at redshifts $z=10$ and 9. Additionally, we explore the impact of these parameters on the observed 21-cm signal power spectrum, focusing on the redshifts within the range of LOFAR 21-cm signal observations ($z\approx 8.5-10$). Despite differences in cosmological and astrophysical parameters, the 21-cm power spectrum at these redshifts agrees with presently observed upper limits. This suggests the need for broader physical parameter spaces for inference modeling to account for all models that agree with observations. However, we also propose stronger constraining power by using a combination of galactic and IGM observables.
@@
pubyear: 2024pagerange: Exploring the effect of different cosmologies on the Epoch of Reionization 21-cm signal with PolarC
@@ 1 Introduction
For this purpose, multiple interferometric low-frequency radio telescopes have been designed, such as PAPER [^1] Precision Array to Probe EoR, http://eor.berkeley.edu, MWA [^2] Murchison Widefield Array, http://www.mwatelescope.org, HERA [^3] Hydrogen Epoch of Reionization Array, https://reionization.org/, LOFAR [^4] Low-Frequency Array, http://www.lofar.org, and the upcoming SKA [^5] Square Kilometre Array, https://www.skao.int/en. While the signal is yet to be detected, upper limits of the 21-cm power spectrum have been obtained and are becoming increasingly tighter (e.g., Mertens et al., 2020; Trott et al., 2020; HERA Collaboration et al., 2023; Acharya et al., 2024c). These have allowed us to rule out some extreme astrophysical models through comparisons with simulations (e.g., Ghara et al., 2020; Mondal et al., 2020; Greig et al., 2021a; Greig et al., 2021b; Abdurashidova et al., 2022). New methods that utilize multi-redshift power spectrum observations (Ghara et al., 2024; Choudhury et al., 2024), wavelet statistics (Hothi et al., 2024), and simulation-based inference (Saxena et al., 2023; Greig et al., 2024) have also been developed to improve the understanding of parameters governing the properties of the IGM.
At the same time, galaxies that formed during the EoR that are responsible for HI ionization have also been studied using hydrodynamical and/or radiative transfer simulations with different prescriptions, such as CROC (Gnedin, 2014; Gnedin & Kaurov, 2014; Esmerian & Gnedin, 2021, 2022, 2024), FIRE (Ma et al., 2018), Technicolor Dawn (Finlator et al., 2018), Sphinx (Rosdahl et al., 2018), CRASH simulations (Eide et al., 2018, 2020; Ma et al., 2021; Kostyuk et al., 2023; Basu et al., 2024), C2Ray simulations (Mellema et al., 2006; Dixon et al., 2016; Giri & Mellema, 2021; Hirling et al., 2024; Giri et al., 2024), CoDa I,II, III (Ocvirk et al., 2016, 2020; Lewis et al., 2022), Astrid (Bird et al., 2022), ColdSIM (Maio et al., 2022; Casavecchia et al., 2024), Thesan (Kannan et al., 2022; Garaldi et al., 2022; Smith et al., 2022; Garaldi et al., 2024), and SPICE (Bhagwat et al., 2024). Although these simulations are ideal tools to investigate the IGM and galactic properties, they are computationally expensive and thus limited in box size and/or in the number of simulations that can be run. Thus, they cannot be used for building a wide range of models for comparison with the 21-cm power spectrum or their upper limits.
@@
The LOFAR EoR Key Science Project (KSP) team has designed Polar, a semi-numeric approach which strikes a balance between speed and complexity of relevant physical processes, by post-processing $N$-body Dark Matter (DM) simulations with Semi-Analytic Models (SAMs) of galaxy formation and evolution, and subsequently applying the 1D radiative transfer (RT) code Grizzly (Ghara et al., 2015; Ghara et al., 2018), an updated version of Bears (Thomas et al., 2009). This allows a fast modeling of the 21-cm signal power spectrum, while at the same time accounting for important galactic properties and RT effects. Consequently, observations of the 21-cm line can also be combined with data obtained in other frequency bands (e.g. from JWST) to jointly constrain IGM and galactic properties.
Here, we investigate the impact of varying cosmological parameters on the 21-cm signal at $z\approx 10.11,9.16$, and 8.3, i.e. the three redshifts targeted by LOFAR (Mertens et al. 2024, in prep). We focus on those parameters whose low and high redshift measurements are in tension, i.e., the Hubble parameter $H_{0}=100h~{}\rm km~{}s^{-1}~{}Mpc^{-1}$ and the matter clustering amplitude $\sigma_{8}$. Our fiducial model is the one of Planck Collaboration et al. (2020), with $h=0.6766$ and $\sigma_{8}=0.8102$, while a higher $h$ value is adopted to explore the low redshift measurement of $h=0.733$ from studies of Cepheid variables in the host galaxies of 42 Type Ia supernovae (Riess et al., 2022). Similarly, a lower $\sigma_{8}$ value of 0.702 is adopted from the low redshift measurement of anisotropic galaxy clustering measurement analysis (Tröster et al., 2020). Lastly, we also adopt a $\sigma_{8}=0.88$ from recent eROSITA results (Ghirardini et al., 2024) to consider both extremes of $\sigma_{8}$ measurements. Motivated by results obtained in the context of the 21-cm signal shown by Giri et al. (2023) and Acharya et al. (2024b), we additionally implement the Fixed & Paired (F&P) approach (Angulo & Pontzen, 2016) to boost effective volumes of the simulations and suppress cosmic variance.
@@ 2 Methodology
Building on M23, we utilize a similar setup by running $N$-body DM simulations, and then post-processing them with the L-Galaxies SAM (Barrera et al., 2023; Henriques et al., 2015, 2020) to model the formation and evolution of galaxies, and with the 1D radiative transfer code Grizzly (Ghara et al., 2015; Ghara et al., 2018) to model the gas ionization and the 21-cm signal from neutral hydrogen. In Sections 2.1 and 2.2, we highlight the key parameters used for setting up the simulations and the analysis done in this work. Further, in Section 2.3 we propose two cases: one where we keep the astrophysical parameters the same for all four cosmological models, and one in which we tune them to match UV luminosity functions (UVLFs) observed with JWST and HST at $z=10$ and 9. We refer to the first and second case as “unconstrained” and “constrained”, respectively.
@@ 2.1 Dark-matter simulations
We use the Gadget-4 code (Springel et al., 2021) for running $N$-body DM simulations with a volume of $(150h^{-1}~{}\rm cMpc)^{3}$ and $2048^{3}$ DM particles. This translates into a DM particle mass of $5\times 10^{7}~{}\rm M_{\odot}$, matching the particle resolution of larger simulations used by the LOFAR EoR KSP team (see for example Giri et al., 2019b, a). The box size is chosen to probe wave-modes of $k\leq 1.0h~{}\rm cMpc^{-1}$, where the best results from observations with LOFAR, HERA, MWA (and eventually SKA) are expected (Koopmans et al., 2015). This condition requires simulations with box sizes $>100h^{-1}~{}\rm cMpc$ (Iliev et al., 2014), although Kaur et al. (2020) suggests box sizes $>175h^{-1}$ cMpc being necessary when accounting for X-ray heating of the IGM. To sidestep this issue, we assume that at our redshifts of interest, the IGM has already been heated above the CMB temperature (see Section 2.2.2). We additionally employ the F&P approach (Angulo & Pontzen, 2016) to mitigate sample variance, and thus run two realisations of each DM simulation. Each pair of realisations has the mode amplitudes fixed to the square root of the initial matter power spectrum, and the phases of the second realisations (B-series) are obtained by mirroring those of the first realisations (A-series). The F&P approach of averaging observables of these two realisations has been shown to boost the statistical precision of the matter power spectrum, bispectrum and halo mass function (Angulo & Pontzen, 2016; Chartier et al., 2021; Maion et al., 2022; Villaescusa-Navarro et al., 2018), Lyman-$\alpha$ power spectra (Anderson et al., 2019), 21-cm signal power spectrum and bispectrum (Giri et al., 2023; Acharya et al., 2024b) and several other quantities derived from simulations (Villaescusa-Navarro et al., 2018; Klypin et al., 2020). Beyond this, we use the same random seed to minimize the effect of randomized initial conditions for both the A and B-series of simulations.
Table 1: The four $N$-body dark matter simulation models considered in this work. From left to right, the model name, the adopted value of $h$ and $\sigma_{8}$, and the reference for the values. All other cosmological parameters are fixed to the Planck Collaboration et al. (2020) values.
@@
For our reference simulation, we assume a “fiducial” $\Lambda$-Cold Dark Matter ($\Lambda$CDM) cosmological model based on Planck Collaboration et al. (2020, specifically the TT,TE,EE+lowE+lensing+BAO case), setting $\Omega_{\Lambda}=0.6889$, $\Omega_{\rm m}=\Omega_{\rm b}+\Omega_{\rm dm}=0.3111$, $\Omega_{\rm b}=0.04897$, $H_{0}=100h~{}\rm km~{}s^{-1}~{}Mpc^{-1}$ with $h=0.6766$, $\sigma_{8}=0.8102$ and $n_{s}=0.9665$, where the symbols have their usual meaning. Further, we consider three additional cosmologies, where we vary $h$ and $\sigma_{8}$, while keeping the other parameters fixed to the above-mentioned values. This is necessary to consider the maximum impact of changing cosmological parameters across the range of observed values. First, we adopt $h=0.7330$ (from Riess et al., 2022), and refer to this as the “$h$ high” model. In addition, we explore two extreme values of $\sigma_{8}$, namely the “$\sigma_{8}$ low” case with $\sigma_{8}=0.702$ (Tröster et al., 2020), and the “$\sigma_{8}$ high” case with $\sigma_{8}=0.88$ (Ghirardini et al., 2024). The details of the four simulation models are listed in Table 1.
The initial conditions for all simulations are generated at $z=199$ with a second-order Lagrangian perturbation theory based on the NgenIC algorithm implemented into Gadget-4, using the same linear theory power spectrum as Hernández-Aguayo et al. (2023). We generate 90 snapshots between $z=20$ and $z=5$, for which we save full snapshot information. However, in order to ensure a maximum step size of 10 Myrs between snapshots used to build halo merger trees with the Friend-Of-Friends (FOF) group finding algorithm (Springel et al., 2001), we perform a finer gridding to generate a total of 156 output time steps for FOF groups. We use a standard linking length of 0.2 times the mean particle spacing and a minimum group size of 64 DM particles (corresponding to a minimum halo size of $\approx 3\times 10^{9}~{}\rm M_{\odot}$). Substructures within halos are identified using the Subfind-HBT algorithm (Han et al., 2018; Springel et al., 2021). Finally, the gravitational softening length is set to 0.025 of the mean particle spacing, i.e. $\approx 1.83h^{-1}~{}\rm ckpc$.
To evaluate the performance of the simulations, in Figure 1 we present the halo mass function (HMF) as $M_{\rm h}~{}\frac{dn_{\rm h}}{dM_{\rm h}}$, where $M_{\rm h}$ and $n_{\rm h}$ are the halo mass and number density respectively, and compare them with fits by Tinker et al. (2008), computed using the Python package hmf (Murray, 2014) at $z=9~{}\rm and~{}10$. As in this study, Tinker et al. (2008) takes into account the non-universality of the HMF by considering the impact of varying the cosmological parameters in the $\Lambda$CDM model. We note that the fiducial and $h$ high models at these redshifts produce very similar HMFs, despite some noticeable differences in their corresponding fits. A deeper exploration indicates that our simulation setup is less sensitive to differences in the Hubble parameter, given the choice of the starting redshift, and that the initial conditions were generated using second-order Lagrangian perturbation theory. A significantly higher starting redshift or third/fourth-order Lagrangian perturbation theory may thus be required to better match the fits. Nevertheless, some differences do exist, especially at the extremities of the HMF, which lead to the differences noted in the subsequent sections.
@@
Figure 1: Halo mass functions for our four models: fiducial (cyan), $h$ high (red), $\sigma_{8}$ low (purple) and $\sigma_{8}$ high (green) at $z=10$ (top panel) and 9 (bottom). The A- and B-series are shown with dashed and dotted lines, respectively. The corresponding fit for each simulation using the Tinker et al. (2008) model is shown with a solid grey line.
@@
Further, we also find a mild under-prediction of the HMF at the low mass end and an over-prediction at the high mass end. This may be due to the choice of the linking length of 0.2, which at significantly high redshifts may lead to overlinking. A lower value of 0.17, as suggested by Watson et al. (2013), may resolve this issue. However, we still use 0.2 for minimizing differences between the setups used for past applications of L-Galaxies (e.g., Henriques et al., 2015, 2020; Barrera et al., 2023, and M23) and our setup. Lastly, we note that the B-series has slightly more massive halos in all models. This is purely because of the choice of the random seed, which happens to lead to some regions of massive halo formation in the B-series where the A-series have voids.

### v2

@@ Abstract
A detection of the 21-cm signal power spectrum from the Epoch of Reionization is imminent, thanks to consistent advancements from telescopes such as LOFAR, MWA, and HERA, along with the development of SKA. In light of this progress, it is crucial to expand the parameter space of simulations used to infer astrophysical properties from this signal. In this work, we explore the role of cosmological parameters such as the Hubble constant $H_{0}$ and the matter clustering amplitude $\sigma_{8}$, whose values as provided by measurements at different redshifts are in tension. We run N-body simulations using Gadget-4, and post-process them with the reionization simulation code Polar, that uses L-Galaxies to include galaxy formation and evolution properties and Grizzly to execute 1-D radiative transfer of ionizing photons in the intergalactic medium (IGM). We compare our results with the latest JWST observations and explore which astrophysical properties for different cosmologies are necessary to match the observed UV luminosity functions at redshifts $z=10$ and 9. Additionally, we explore the impact of these parameters on the observed 21-cm signal power spectrum upper limits, focusing on the redshifts within the range of LOFAR 21-cm signal observations ($z\approx 8.5-10$). Despite differences in cosmological and astrophysical parameters, our models cannot be ruled out by the current upper limits. This suggests the need for broader physical parameter spaces for inference modeling to account for all models that agree with observations. However, we also propose stronger constraining power by using a combination of galactic and IGM observables.
@@
pubyear: 2024pagerange: Exploring the effect of different cosmologies on the Epoch of Reionization 21-cm signal with PolarC.2
@@ 1 Introduction
For this purpose, multiple interferometric low-frequency radio telescopes have been designed, such as PAPER [^1] Precision Array to Probe EoR, http://eor.berkeley.edu, MWA [^2] Murchison Widefield Array, http://www.mwatelescope.org, HERA [^3] Hydrogen Epoch of Reionization Array, https://reionization.org/, LOFAR [^4] Low-Frequency Array, http://www.lofar.org, and the upcoming SKA [^5] Square Kilometre Array, https://www.skao.int/en. While the signal is yet to be detected, upper limits of the 21-cm power spectrum have been obtained and are becoming increasingly tighter (e.g., Mertens et al., 2020; Trott et al., 2020; HERA Collaboration et al., 2023; Acharya et al., 2024c). These have allowed us to rule out some extreme astrophysical models through comparisons with simulations (e.g., Ghara et al., 2020; Mondal et al., 2020; Greig et al., 2021a; Greig et al., 2021b; Abdurashidova et al., 2022). New methods that utilise multi-redshift power spectrum observations (Ghara et al., 2024; Choudhury et al., 2024), wavelet statistics (Hothi et al., 2024), and simulation-based inference (Saxena et al., 2023; Greig et al., 2024) have also been developed to improve the understanding of parameters governing the properties of the IGM.
At the same time, galaxies that formed during the EoR that are responsible for HI ionization have also been studied using hydrodynamical and/or radiative transfer simulations with different prescriptions, such as CROC (Gnedin, 2014; Gnedin & Kaurov, 2014; Esmerian & Gnedin, 2021, 2022, 2024), FIRE (Ma et al., 2018), Technicolor Dawn (Finlator et al., 2018), Sphinx (Rosdahl et al., 2018), CRASH simulations (Eide et al., 2018, 2020; Ma et al., 2021; Kostyuk et al., 2023; Basu et al., 2024), C2Ray simulations (Mellema et al., 2006; Dixon et al., 2016; Giri & Mellema, 2021; Hirling et al., 2024; Giri et al., 2024), CoDa I,II, III (Ocvirk et al., 2016, 2020; Lewis et al., 2022), Astrid (Bird et al., 2022), ColdSIM (Maio et al., 2022; Maio & Viel, 2023; Casavecchia et al., 2024), Thesan (Kannan et al., 2022; Garaldi et al., 2022; Smith et al., 2022; Garaldi et al., 2024), and SPICE (Bhagwat et al., 2024). Although these simulations are ideal tools to investigate the IGM and galactic properties, they are computationally expensive and thus limited in box size and/or in the number of simulations that can be run. Thus, they cannot be used for building a wide range of models for comparison with the 21-cm power spectrum or their upper limits.
@@
The LOFAR EoR Key Science Project (KSP) team has designed Polar, a semi-numeric approach that strikes a balance between speed and complexity of relevant physical processes, by post-processing $N$-body Dark Matter (DM) simulations with Semi-Analytic Models (SAMs) of galaxy formation and evolution, and subsequently applying the 1D radiative transfer (RT) code Grizzly (Ghara et al., 2015; Ghara et al., 2018), an updated version of Bears (Thomas et al., 2009). This allows a fast modelling of the 21-cm signal power spectrum, while at the same time accounting for important galactic properties and RT effects. The 21-cm signal can be used to derive physical properties of the Universe (McQuinn et al., 2006), but as it primarily comes from the IGM, it provides only a partial picture. Thus, instead, we combine observations of the 21-cm line with data obtained in other frequency bands (e.g. from JWST) to jointly constrain IGM and galactic properties. This has been explored in the past, for example, by using cross-power spectra of galaxy clustering observations (Lidz et al., 2009; Vrbanec et al., 2016; Berti et al., 2024), or intensity mapping (Mao et al., 2008; La Plante et al., 2023; Berti et al., 2024) at different redshifts. Eventually, the goal of such studies would be to jointly constrain the physical properties of the Universe from multiple observables of galaxies and the IGM.
[…]
Our fiducial model is the one of Planck Collaboration et al. (2020), with $h=0.6766$ and $\sigma_{8}=0.8102$, while a higher $h$ value is adopted to explore the low redshift measurement of $h=0.733$ from studies of Cepheid variables in the host galaxies of 42 Type Ia supernovae (Riess et al., 2022). Similarly, a lower $\sigma_{8}$ value of 0.702 is adopted from the low-redshift measurement of anisotropic galaxy clustering measurement analysis (Tröster et al., 2020). Lastly, we also adopt a $\sigma_{8}=0.88$ from recent eROSITA results (Ghirardini et al., 2024) to consider both extremes of $\sigma_{8}$ measurements. Motivated by results obtained in the context of the 21-cm signal shown by Giri et al. (2023) and Acharya et al. (2024b), we additionally implement the Fixed & Paired (F&P) approach (Angulo & Pontzen, 2016) to boost effective volumes of the simulations and suppress cosmic variance.
@@ 2 Methodology
Building on M23, we utilise a similar setup by running $N$-body DM simulations, and then post-processing them with the L-Galaxies SAM (Barrera et al., 2023; Henriques et al., 2015, 2020) to model the formation and evolution of galaxies, and with the 1D radiative transfer code Grizzly (Ghara et al., 2015; Ghara et al., 2018) to model the gas ionization and the 21-cm signal from neutral hydrogen. In Sections 2.1 and 2.2, we highlight the key parameters used for setting up the simulations and the analysis done in this work. Further, in Section 2.3 we propose two cases: one where we keep the astrophysical parameters the same for all four cosmological models, and one in which we tune them to match UV luminosity functions (UVLFs) observed with JWST and HST at $z=10$ and 9. We refer to the first and second case as “unconstrained” and “constrained”, respectively.
@@ 2.1 Dark-matter simulations
We use the Gadget-4 code (Springel et al., 2021) for running $N$-body DM simulations with a volume of $(150h^{-1}~\rm cMpc)^{3}$ and $2048^{3}$ DM particles. This translates into a DM particle mass of $5\times 10^{7}~\rm M_{\odot}$, matching the particle resolution of larger simulations used by the LOFAR EoR KSP team (see for example Giri et al., 2019b, a). The box size is chosen to probe wave-modes of $k\leq 1.0h~\rm cMpc^{-1}$, where the best results from observations with LOFAR, HERA, MWA (and eventually SKA) are expected (Koopmans et al., 2015). This condition requires simulations with box sizes $100h^{-1}~\rm cMpc$ (Iliev et al., 2014), although Kaur et al. (2020) suggests box sizes $175h^{-1}$ cMpc being necessary when accounting for X-ray heating of the IGM. To sidestep this issue, we assume that at our redshifts of interest, the IGM has already been heated above the CMB temperature (see Section 2.2.2). We additionally employ the F&P approach (Angulo & Pontzen, 2016) to mitigate sample variance, and thus run two realisations of each DM simulation. Each pair of realisations has the mode amplitudes fixed to the square root of the initial matter power spectrum, and the phases of the second realisations (B-series) are obtained by mirroring those of the first realisations (A-series). This is in contrast to the traditional method of generating initial conditions, where the mode amplitudes are sampled from a Gaussian distribution centred on the expectation value, which is given by the initial matter power spectrum. Taking the F&P approach suppresses the variance-induced fluctuations (which in traditional initial conditions mainly affect the large-scale modes, where the sampling of the power spectrum amplitude is scarce and therefore more susceptible to deviations from its expectation value) up to the fourth perturbative order (Angulo & Pontzen, 2016). This means that while small-scale properties, such as individual halos, may not show any noticeable difference, averaged large-scale statistics of an F&P pair reproduce very closely the ensemble averages of traditional simulations. The F&P approach of averaging observables of these two realisations has been shown to boost the statistical precision of the matter power spectrum, bispectrum and halo mass function (Angulo & Pontzen, 2016; Chartier et al., 2021; Maion et al., 2022; Villaescusa-Navarro et al., 2018), Lyman-$\alpha$ power spectra (Anderson et al., 2019), 21-cm signal power spectrum and bispectrum (Giri et al., 2023; Acharya et al., 2024b) and several other quantities derived from simulations (Villaescusa-Navarro et al., 2018; Klypin et al., 2020). Beyond this, we use the same random seed to minimise the effect of randomized initial conditions for both the A and B-series of simulations.
Table 1: The four NN-body dark matter simulation models considered in this work. From left to right, the model name, the adopted value of $h$ and $\sigma_{8}$, and the reference for the values. All other cosmological parameters are fixed to the Planck Collaboration et al. (2020) values.
@@
[…]
@@
Figure 1: Halo mass functions (HMFs) for our four models: fiducial (cyan), $h$ high (red), $\sigma_{8}$ low (purple) and $\sigma_{8}$ high (green) at $z=10$ (top panel) and 9 (bottom). The A- and B-series are shown with dashed and dotted lines, respectively. The corresponding fit for each simulation using the Tinker et al. (2008) model is shown with solid lines of the corresponding colours.
The initial conditions for all simulations are generated at $z=199$ with a second-order Lagrangian perturbation theory based on the NgenIC algorithm implemented into Gadget-4, using the same linear theory power spectrum as Hernández-Aguayo et al. (2023). We generate 90 snapshots between $z=20$ and $z=5$, for which we save the full snapshot information, i.e., the positions and velocities of all the dark matter particles. However, in order to ensure a maximum step size of 10 Myrs between snapshots used to build halo merger trees with the Friend-Of-Friends (FOF) group finding algorithm (Springel et al., 2001), we perform a finer gridding to generate a total of 156 output time steps for FOF groups. This is necessary because the dynamical time of a typical galactic disc at high redshift is of the order of tens of Myrs (Poole et al., 2016; Balu et al., 2023), which also is the lifetime of massive stars (Barkana & Loeb, 2001; Schaerer, 2002; Heger & Woosley, 2002; Bromm & Larson, 2004). Thus, a longer time step would miss how reionization progresses when post-processed with SAMs like L-Galaxies and radiative transfer codes like Grizzly. We use a standard linking length of 0.2 times the mean particle spacing and a minimum group size of 64 DM particles (corresponding to a minimum halo size of $\approx 3\times 10^{9}~\rm M_{\odot}$). Substructures within halos are identified using the Subfind-HBT algorithm (Han et al., 2018; Springel et al., 2021). Finally, the gravitational softening length is set to 0.025 of the mean particle spacing, i.e. $\approx 1.83h^{-1}~\rm ckpc$.
[…]
@@
Further, we also find a mild under-prediction of the HMF at the low mass end and an over-prediction at the high mass end. This may be due to the choice of the linking length of 0.2, which at significantly high redshifts may lead to overlinking. A lower value of 0.17, as suggested by Watson et al. (2013), may resolve this issue. However, we still use 0.2 for minimizing differences between the setups used for past applications of L-Galaxies (e.g., Henriques et al., 2015, 2020; Barrera et al., 2023, and M23) and our setup. Lastly, we note that the B-series has slightly more massive halos in all models. This is purely because of the choice of the random seed; the same seed that leads to voids in the A-series produces regions of massive halo formation in the B-series due to the F&P method described above.